A piece of copper weighing 300g is heated to 100°c and is then quickly transferred into a copper calorimeter of mass who containing 10g containing 100g of liquid with unknown specific heat capacity. If the final temperature of the mixture is 40°c calculate the specific heat capacity of the liquid (room temperature is 10°c)

(Specific heat capacity of copper is 400jkg/1


1
Expert's answer
2021-04-13T06:40:53-0400

Writing down the termal balance equation, obtain:


"Q_p =Q_c+Q_l"

which means that the extra termal energy of the piece of copper "Q_p" was spent to increase the termal energies of calorimeter "Q_c" and liquid "Q_l". Substituting the expressions for energies, obtain:


"c_cm_p(T_1-T_0) = c_lm_l(T_0-T_2) + c_cm_c(T_0-T_2)"

where "c_c =400\frac{J}{kg\cdot \degree C}" is the specific heat capacity of copper, "c_l" is the unknown specific heat capacity of liquid, "m_p = 300g = 0.3kg, m_c = 10g = 0.01kg, m_l = 100g = 0.1kg" are the masses of the piece, calorimeter and liquid respectively, "T_0 = 40\degree C" is the final temperature, "T_1 = 100\degree C" is the initial temperature of the piece, and "T_2 = 10\degree C" is the initial temperature of the calorimeter and liquid.

Expressing "c_l", find:


"c_l = \dfrac{c_cm_p(T_1-T_0) - c_cm_c(T_0-T_2)}{ m_l(T_0-T_2)} = \dfrac{c_cm_p(T_1-T_0)}{m_l(T_0-T_2)}-\dfrac{c_cm_c}{m_l}\\
c_l =\dfrac{400\cdot 0.3\cdot (100-40)}{0.1\cdot (40-10)} - \dfrac{400\cdot 0.01}{0.1} =2360\dfrac{J}{kg\cdot \degree C}"

Answer. "2360\dfrac{J}{kg\cdot \degree C}".